# SLoRD: Structural Low-Rank Descriptors for Shape Consistency in Vertebrae Segmentation

### v2

@@ Abstract
Automatic and precise multi-class vertebrae segmentation from CT images is crucial for various clinical applications. However, due to a lack of explicit consistency constraints, existing methods especially for single-stage methods, still suffer from the challenge of intra-vertebrae segmentation inconsistency, which refers to multiple label predictions inside a singular vertebra. For multi-stage methods, vertebrae detection serving as the first step, tends to be affected by the pathology and metal implants. Thus, imprecise detections cause biased patches before segmentation, which then leads to inaccurate contour delineation and inconsistent segmentation. In our work, we intend to label individual and complete binary masks to address that challenge. Specifically, a contour generation network is proposed based on Structural Low-Rank Descriptors for shape consistency, termed SLoRD. For a structural representation of vertebral contours, we adopt the spherical coordinate system and devise the spherical centroid to calculate contour descriptors. Due to vertebrae’s similar appearances, basic contour descriptors can be acquired to restore original contours. Therefore, SLoRD leverages these contour priors and explicit shape constraints to facilitate regressed contour points close to vertebral surfaces. Quantitative and qualitative evaluations on VerSe 2019 and 2020 demonstrate the superior performance of our framework over other single-stage and multi-stage state-of-the-art (SOTA) methods. Further, SLoRD is a plug-and-play framework to refine the segmentation inconsistency existing in coarse predictions from other approaches. Source codes are available https://github.com/AlexYouXin/SLoRD-VerSe.
@@
Vertebrae segmentation, Inconsistency, Contour descriptors, Refinement
@@ 1 Introduction
Automatic multi-class vertebrae segmentation from CT images plays a significant role in spinal disease treatment, including preoperative diagnosis, surgical guidance, and postoperative assessment [22, 5, 21]. Moreover, the shape analysis of vertebrae serve as an important anatomical reference for other anatomies in clinical practices [9, 28]. According to the anatomical prior of vertebral structures, it is emerging to preserve the consistency of label predictions inside a singular vertebra [24]. However, as illustrated in Fig.1, for single-stage and multi-stage methods, predicted masks for the specific vertebra might contain inconsistent label predictions, and partial voxel predictions in adjacent vertebrae reveal the same label. That phenomenon has been recognized as intra-vertebrae label inconsistency [45], which poses an enormous challenge for the task of precise vertebrae segmentation.
@@
Specifically, earlier methods solved vertebrae segmentation from volumetric CT images by employing statistical models [20], active shape models [14], level sets [23], Markov random field [17, 18], graph cuts [3] and machine learning theories [10, 19]. However, those approaches fail to perform well on unseen data due to insufficient generalization ability. Current methods based on deep learning techniques demonstrate a more promising performance for vertebrae image segmentation even with arbitrary field-of-views (FOVs). One category of existing works [11, 15, 45] adopted single-stage convolutional or Transformer-based networks for automatic vertebrae segmentation and identification. However, these models suffer from drawbacks of similar appearances between adjacent vertebrae and the patch-based inference paradigm [22], resulting in intra-vertebrae label inconsistency as revealed in Fig.1(a & b). The other methods [27, 26, 38, 31] designed a multi-stage pipeline, which can boost the segmentation performance by reducing the disturbance from neighboring contexts based on the localize-then-segment concept. However, a biased local patch from the detection network due to the inaccurate center regression will enforce inconsistency inside vertebrae, particularly for pathological vertebrae with abnormal shapes or metal implants as illustrated by Fig.1(d & e). Also, this limitation can be referred to according to the qualitative analysis of experimental results in [38].
@@
Essentially, all the above methods fail to explicitly impose both contour precision and intra-vertebrae voxel consistency constraints synchronously, resulting in segmentation inconsistency. To resolve this limitation, we attempt to label individual and complete binary masks to achieve multi-label vertebrae segmentation. Due to the connection between superior and inferior articular processes [5], the main challenge is transformed to how to divide joint binary vertebral masks while preserving original contours to the greatest extent [13].
@@
Figure 2: Spherical coordinate system for 3D vertebral contour descriptions. x, y, and z axis represent voxel orientations from Right/Anterior/Inferior to Left/Posterior/Superior. $\rho$: the radial distance away from the spherical center, $\phi$: the azimuth angle between the radial vector and $z^{+}$, $\theta$: the polar angle between $x^{+}$ and the projected radial vector in the XOY plane.
@@
Thus, we devise an effective contour generation network based on Structural Low-Rank Descriptors for shape consistency, termed SLoRD. For the contour representation, the complete point cloud of vertebral boundaries shows an unorganized arrangement with various point numbers. Intuited by [43], we propose a novel way of contour representations based on the spherical coordinate system as shown in Fig.2. Under that circumstance, structural contour descriptors are efficiently calculated by recording the radial distance. Additionally, spherical centroids are proposed to enhance the precision of contour representations, serving as the spherical center. Then, vertebrae show similar 3D shapes according to anatomical structures, lumbar and thoracic vertebrae in particular. Hence, a low-dimension contour space can be extracted to represent all vertebrae data. Via a linear combination between low-rank basic descriptors in the contour space, instance masks of vertebrae can be approximately restored. Further, vertebral labels provided by coarse predictions facilitate the consistent generation of semantic masks. Ultimately, the whole pipeline is incorporated into SLoRD to boost the framework’s generalization abilities on unseen datasets. And the explicit shape constraint and basic contour descriptors are exerted to promote the regression precision of contours. SLoRD serves as a plug-and-play framework, which can repair the segmentation inconsistency from coarse predictions.
@@ 2 Related work
For single-stage pipelines, Zheng et al. employed nnUNet [15] to perform multi-class vertebrae segmentation. Chang et al. [6] adopted stacked graph convolutional networks by featuring a pre-defined adjacency matrix to encode structural information. However, such approaches fail to model global dependency due to the intrinsic property of convolution [12]. Thus, Verteformer [45] was devised by fusing local details and global contexts with two parallel Transformers. And an edge detection block was proposed to refine intra-vertebrae boundaries. Nevertheless, it is still plagued by similar shapes between adjacent vertebrae and the patch-based inference paradigm, leading to intra-vertebrae segmentation inconsistency. Lessmann et al. [22] used fully convolutional networks by iteratively analyzing image patches for automatic vertebrae segmentation and identification. However, this method relies on the perfect prediction for the first label and hardly recovers from failures in the top vertebra.
@@ 3.1 Spherical-based Contour Descriptors
Although the number of contour points varies across different vertebrae, a fixed number of sampling points is required for contour representations. However, it is quite tricky to sample contour points uniformly while maintaining the potential to restore the original contour. Thus, we incorporate the idea from PolarMask [43, 30], in which each contour point is encoded with the polar coordinate.
Figure 4: The two-stage framework. (a) Coarse predictions by arbitrary segmentation networks. (b) The generation of Gaussian positional prompts to boost the center localization precision. (c) Based on the combination of basic contour descriptors, SLoRD aims to regress contour points close to vertebral boundaries via explicit contour and shape regularizations.
Given the difficulty in extending the polar system to contour representations of 3D objects like vertebrae, we formulate contours with the spherical coordinate system as shown in Fig.2. Conditioning the inner centroid as the spherical center, we describe the boundary using spherical coordinates $(\rho_{{}_{\mathrm{ij}}},\theta_{i},\phi_{j}),i\mbox{=}1,...,I;j\mbox{=}1,...,J$. Here $\rho$ represents the length of the radial vector from the spherical center to the contour point, $\phi$ is the azimuth angle between the radial vector and $z^{+}$, $\theta$ means the polar angle between $x^{+}$ and the projected radial vector in the XOY plane. Angular coordinates $\theta_{i}$ and $\phi_{j}$ are sampled uniformly with a sampling interval $s$, with $I$ equal to $360/s$, $J$ equal to $(180/s\mbox{+}1)$. Thus, given the sampling angle grid, only radial coordinates are recorded to represent the contour:
@@
Here $\bm{\rho}$ refers to the contour descriptor, with the dimension $N$ equal to $I\times J$. Considering vertebral anatomies with the non-convex shape property, the spherical system chooses the maximum distance at the same angle to strike a balance between precision and computational cost. Since the shapes of most vertebrae are well structured and resembling to each other, low-rank contour descriptors can be exploited in a data-driven manner via Singular Value Decomposition (SVD) [30]. Detailedly, a non-convex contour matrix $\mathcal{M}=[\bm{\rho}_{1},\cdot\cdot\cdot,\bm{\rho}_{L}]$ is constructed from L vertebrae in the training data. Then we perform SVD on matrix $\mathcal{M}$:
@@
Another relevant research DeepSSM [4], as a shape model based on principal component analysis (PCA), aims to learn the functional mapping from images to low-dimensional latent codes. Then the model achieves discrete contour representations by applying the linear projection on these codes with the PCA basis and mean shape. In comparison, SVD bears the best low-rank approximation properties, and extracted basic descriptors are more robust to fit various anatomical shapes than PCA [40]. Also, the sampling strategy in [4] for the contour point cloud struggles to perform well for pathological vertebral structures with more diverse shapes and topology.
@@ 3.2 Spherical Centroid
Although the encoding scheme for vertebral contours is efficient, it is hard to determine the spherical center, which will significantly affect the restoration performance of vertebral shapes. Since the azimuth angle $\phi$ and polar angle $\theta$ are both sampled at equal intervals, a spherical center with a biased location will result in fewer sampling contour points for the boundary region far from it. Consequently, that region suffers from a contour under-representation. Hence, it has been a crucial topic on how to designate the spherical center.
A simple configuration is each vertebral centroid, averaged on both the vertebral body and spinous process. However, the centroid is acquired ignoring the optimization objective of enforcing contour points close to the boundary $\mathcal{B}$. Thus, to achieve a finer contour representation for the anatomical structure of vertebrae, a distance-based center should be marked by minimizing the average Euclidean distance between this center and the whole boundary surface. We formulate the distance-based center called spherical centroid $C_{s}$ based on Eq (4):
@@
where $\mathcal{F}(\cdot)$ represents the morphological dilation operation, $M$ refers to the segmentation mask of each vertebral volume, the boundary map $\mathcal{B}$ is the difference set between the dilated mask $\mathcal{F}(M)$ and the mask itself $M$. Besides, $d(\cdot)$ means the Euclidean distance. The first item in Eq (4) refers to the average distance from each boundary point $b$ to the center $c\in\mathcal{P}$. Concerning the non-convex shape prior of the spinous process, we attempt to push the spherical centroid closer to it. That will enlarge the value range of $\theta$ and $\phi$, then enhance the shape representations of the spinous process. That is why the second regularization term is incorporated, in which $\delta$ is the boundary point $j\in\mathcal{B}$ with the largest y-dimensional index $j_{y}$. $\lambda$ is a scaling parameter, empirically set as $0.005$.
@@ 3.3 Contour Generation with Explicit Contour Priors
Center decoder: Deep features from the encoder are transformed into regressed spherical centers via combinative operations of convolution, batch normalization, average pooling. The pooling process projects the feature dimension as 3, representing coordinates of the regressed center $\hat{C}=(\hat{c}_{x},\hat{c}_{y},\hat{c}_{z})$. The ground truth of center $C_{s}$ can be attained by Eq (4). And we employ the $L_{2}$ norm to regularize regressed centers $\hat{C}$, with the center loss $\mathcal{L}_{center}$ equal to $||\,C_{s}-\hat{C}\,||^{2}$. Concerning the limitation that some methods wrongly detect centers of specific pathological vertebrae, we aim to introduce Gaussian positional prompts as additional inputs for positional guidance. As shown in Fig.4(b), the positional prompt is defined as a Gaussian distribution $\mathcal{G}$ based on the follow formula:
@@
Specifically, the mean value $\bm{\mu}=(\mu_{x},\mu_{y},\mu_{z})$ is set as the centroid based on the coarse prediction. For defining the variance $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$, we aim to cover the whole shape distribution using the given $\mathcal{G}$. Thus, as suggested by the two-sigma rule in mathematical statistics, the length of $4\bm{\sigma}$ should encompass the 3-dimension length $\bm{m}=(m_{x},m_{y},m_{z})$ of each instance mask. Further, a threshold $\bm{\bar{m}}$ is introduced to ensure a wide enough region with attention weights decaying from the center $\bm{\mu}$ to the vertebral boundary. The center decoder will modify coarse centroids $\bm{\mu}$ as refined spherical centroids via $\mathcal{L}_{center}$.
@@
Loss term: To reduce computational costs, we uniformly sample $\tfrac{1}{3}$ boundary points from $\mathcal{B}$ to calculate the contour loss. Besides, an additional binary mask supervision $\mathcal{L}_{mask}$ (cross-entropy + Dice loss) will exert implicit shape information. Thus, the total loss is defined as follows:
@@
3.4 Inference
In the inference phase, the patch to be refined in SLoRD is cropped along each vertebral center from coarse predictions as shown in Fig.4(a). The SLoRD framework will first refine three spherical centers corresponding to sequential vertebrae according to the prediction of the center decoder. Then the coefficient branch will output $3\times N$ (N refers to the dimension of $\bm{\rho}$) contour points surrounding three vertebral surfaces with different labels as revealed in Fig.5. After that, Marching Cube [25] is implemented to generate triangle meshes, which are further voxelized into a filled volume [33]. After that, we perform the binarization on the coarse mask and multiply it with the filled volume to realize a refined mask with good label consistency and shape restoration via the auxiliary connected component analysis.

### v3

@@ Abstract
Automatic and precise multi-class vertebrae segmentation from CT images is crucial for various clinical applications. However, due to similar appearances between adjacent vertebrae and the existence of various pathologies, existing single-stage and multi-stage methods suffer from imprecise vertebrae segmentation. Essentially, these methods fail to explicitly impose both contour precision and intra-vertebrae voxel consistency constraints synchronously, resulting in the intra-vertebrae segmentation inconsistency, which refers to multiple label predictions inside a singular vertebra. In this work, we intend to label complete binary masks with sequential indices to address that challenge. Specifically, a contour generation network is proposed based on Structural Low-Rank Descriptors for shape consistency, termed SLoRD. For a structural representation of vertebral contours, we adopt the spherical coordinate system and devise the spherical centroid to calculate contour descriptors. Due to vertebrae’s similar appearances, basic contour descriptors can be acquired offline to restore original contours. Therefore, SLoRD leverages these contour priors and explicit shape constraints to facilitate regressed contour points close to vertebral surfaces. Quantitative and qualitative evaluations on VerSe 2019 and 2020 demonstrate the superior performance of our framework over other single-stage and multi-stage state-of-the-art (SOTA) methods. Further, SLoRD is a plug-and-play framework to refine the segmentation inconsistency existing in coarse predictions from other approaches. Source codes are available https://github.com/AlexYouXin/SLoRD-VerSe.
@@
Vertebrae segmentation, Inconsistency, Contour descriptors, Refinement.
@@ 1 Introduction
Automatic multi-class vertebrae segmentation from CT images with arbitrary field-of-views (FOVs) plays a significant role in spinal disease treatment, including preoperative diagnosis, surgical guidance, and postoperative assessment [1, 2, 3]. Moreover, the shape analysis of vertebrae serve as an important anatomical reference for other anatomies in clinical practices [4, 5]. According to the anatomical prior of vertebral structures, it is emerging to preserve the consistency of label predictions inside a singular vertebra [1]. However, due to similar appearances between adjacent vertebrae and the existence of various pathologies [6, 7, 8, 9], single-stage and multi-stage methods are prone to generate inconsistent mask predictions, which contain inconsistent label predictions for a singular vertebra as illustrated in Fig.1. That phenomenon can be recognized as intra-vertebrae label inconsistency [10], which poses an enormous obstacle for the task of precise vertebrae segmentation.
@@
Specifically, earlier methods solved vertebrae segmentation from volumetric CT images by employing statistical models [11], active shape models [12], level sets [13], Markov random field [14, 15], graph cuts [16] and machine learning theories [17, 18]. However, those approaches fail to perform well on unseen data due to insufficient generalization ability. Current methods based on deep learning techniques demonstrate a more promising performance for vertebrae image segmentation even with arbitrary FOVs. One category of existing works [19, 20, 10] adopted single-stage convolutional or Transformer-based networks for automatic vertebrae segmentation and identification. However, these models suffer from drawbacks of similar appearances between adjacent vertebrae and the patch-based inference paradigm [1], resulting in intra-vertebrae label inconsistency as revealed in Fig.1(a & b). The other methods [21, 22, 7, 23] designed a multi-stage pipeline, which can boost the segmentation performance by reducing the disturbance from neighboring contexts based on the localize-then-segment concept. However, a biased local patch from the detection network due to the inaccurate center regression will enforce inconsistency inside vertebrae, particularly for pathological vertebrae with abnormal shapes or metal implant artifacts as illustrated by Fig.1(d & e). Also, this limitation can be referred to according to the qualitative analysis of experimental results in [7].
@@
Essentially, all the above methods fail to explicitly impose both contour precision and intra-vertebrae voxel consistency constraints synchronously, resulting in segmentation inconsistency. To resolve this limitation, we attempt to label complete binary masks with sequential indices to achieve multi-label vertebrae segmentation. Due to the connection between superior and inferior articular processes [2], the main challenge is transformed to how to divide joint binary vertebral masks while preserving original contours to the greatest extent [24].
@@
Figure 2: Spherical coordinate system for contour representations of 3D vertebrae. (a). red point: the spherical center, green point: the contour point, blue line: the radial vector. (b). x, y, and z axis adopt the orientations shown in the figure. $\rho$: the radial distance away from the spherical center, $\phi$: the azimuth angle between the radial vector and $z^{+}$, $\theta$: the polar angle between $x^{+}$ and the projected radial vector in the XOY plane.
@@
Thus, we devise an effective contour generation network based on Structural Low-Rank Descriptors for shape consistency, termed SLoRD. For the contour representation, the complete point cloud of vertebral boundaries shows an unorganized arrangement with various point numbers. Intuited by [25], we propose a novel way of contour representations based on the spherical coordinate system as shown in Fig.2. Under that circumstance, structural contour descriptors are efficiently calculated by recording the radial distance. Additionally, spherical centroids are proposed to enhance the precision of contour representations, serving as the spherical center. Then, vertebrae show similar 3D shapes according to anatomical structures, lumbar and thoracic vertebrae in particular. Hence, low-rank contour descriptors can be extracted from the training dataset in an offline manner. Via a linear combination between low-rank basic descriptors in the contour space, instance masks of vertebrae can be approximately restored. Further, vertebral labels provided by coarse predictions facilitate the consistent generation of semantic masks. Ultimately, the whole pipeline is incorporated into SLoRD to boost the framework’s generalization abilities on unseen datasets. And the explicit shape constraint and basic contour descriptors are exerted to promote the regression precision of contours. SLoRD serves as a plug-and-play framework, which can repair the segmentation inconsistency from coarse predictions.
@@ 2 Related work
For single-stage pipelines, Zheng et al. employed nnUNet [20] to perform multi-class vertebrae segmentation. Chang et al. [33] adopted stacked graph convolutional networks by featuring a pre-defined adjacency matrix to encode structural information. However, such approaches fail to model global dependency due to the intrinsic property of convolution [34]. Thus, Verteformer [10] was devised by fusing local details and global contexts with two parallel Transformers. And an edge detection block was proposed to refine intra-vertebrae boundaries. Nevertheless, it is still plagued by similar shapes between adjacent vertebrae and the patch-based inference paradigm, leading to intra-vertebrae segmentation inconsistency. Lessmann et al. [1] used fully convolutional networks by iteratively analyzing image patches for automatic vertebrae segmentation and identification. However, this method relies on the perfect localization of the topmost vertebra using a sliding window fashion, thus is prone to sequential poor segmentation without a precise detection of the topmost vertebra.
@@ 3.1 Spherical-based Contour Descriptors
Although the number of contour points varies across different vertebrae, a fixed number of sampling points is required for contour representations due to the predefined network. However, it is quite tricky to sample contour points uniformly while maintaining the potential to restore the original contour. Thus, we incorporate the idea from PolarMask [25, 37], in which each contour point is encoded with the polar coordinate.
Figure 4: The two-stage framework. (a) The first stage adopts arbitrary segmentation networks to output inconsistent predictions, which will be refined with intra-vertebrae segmentation consistency by SLoRD in a sliding-window fashion. (b) Gaussian positional prompts are acquired based on the position of coarse spherical centers and the size of instance masks. Instance-wise positional priors are beneficial to promote the spherical centroid localization by SLoRD. (c) SLoRD consists of the spherical center decoder, the coefficient decoder, and the binary mask decoder. Specifically, the center decoder will regress spherical centroids with better precision. The coefficient decoder will generate linear coefficients, which are transformed into contour points under the interaction with low-rank contour descriptors. The mask decoder is aimed for precise shape representations of vertebrae, which boost the precision of generated contours. SLoRD aims to regress contour points close to vertebral boundaries via explicit contour and shape regularizations.
Given the difficulty in extending the polar system to contour representations of 3D objects like vertebrae, we formulate contours with the spherical coordinate system as shown in Fig.2. Conditioning the inner centroid as the spherical center, we describe the boundary using spherical coordinates $(\rho,\theta,\phi)$. Here $\rho$ represents the length of the radial vector from the spherical center to the contour point, $\phi$ is the azimuth angle between the radial vector and $z^{+}$, $\theta$ means the polar angle between $x^{+}$ and the projected radial vector in the XOY plane. Angular coordinates $\theta_{i}$ and $\phi_{j}$ are sampled uniformly with a sampling interval $s$, with $I$ equal to $360/s$, $J$ equal to $(180/s\mbox{+}1)$. Thus, given the sampling angle grid, only radial coordinates are recorded to represent the contour:
@@
$\rho_{{}_{\mathrm{ij}}}=\rho(\theta_{i},\phi_{j}),\,\,i=1,...,I;\;j=1,...,J$ (2)
here $\bm{\rho}$ refers to the contour descriptor, with the dimension $N$ equal to $I\times J$. $\rho_{ij}$ means the radial length from the spherical center to the specific contour point with the polar angle and azimuth angle equal to $\theta_{i}$ and $\phi_{i}$. Considering vertebral anatomies with the non-convex shape property, the spherical system chooses the maximum distance at the same angle to strike a balance between precision and computational cost. Since the shapes of most vertebrae are well structured and resembling to each other, low-rank contour descriptors can be exploited in a data-driven manner via Singular Value Decomposition (SVD) [37]. Detailedly, a non-convex contour matrix $\mathcal{M}=[\bm{\rho}_{1},\cdot\cdot\cdot,\bm{\rho}_{L}]$ is constructed from L vertebrae in the training data. Then we perform SVD on matrix $\mathcal{M}$:
@@
Another contour-based DeepSSM [38], as a shape model based on principal component analysis (PCA), aims to learn the functional mapping from images to low-dimensional latent codes $\mathcal{C}_{l}$. Then the model achieves discrete contour representations $\mathcal{R}_{c}$ by applying the linear projection on these codes with the PCA basis $\mu$ and mean shape $\nu$. This process can be described as follows:
$\mathcal{R}_{c}=\mu\,\mathcal{C}_{l}+\nu$ (5)
In contrast, SVD bears the best low-rank approximation properties, and extracted basic descriptors are more robust to fit various anatomical shapes than PCA [39]. Also, the sampling strategy in [38] for the contour point cloud struggles to perform well for pathological vertebral structures with more diverse shapes and topology.
@@ 3.2 Spherical Centroid
Since the azimuth angle $\phi$ and polar angle $\theta$ are both sampled at equal intervals, a spherical center with a biased location will result in sparsely sampled contour points for the boundary region far from the center. Consequently, that region suffers from a contour under-representation, which further affects the restoration of vertebral shapes. Hence, it is crucial to designate the spherical center adaptive to the specific shape information of vertebrae.
A simple and straightforward configuration is the vertebral centroid, averaged on both the vertebral body and spinous process. However, the centroid is acquired ignoring the optimization objective of enforcing contour points close to the boundary $\mathcal{B}$. Specifically, in the process of constructing the contour matrix $\mathcal{M}$, angles corresponding to boundary points are represented as floating-point numbers. To improve computational and storage efficiency, these angles are approximated to integers. Consequently, sampling contour points based on an integer-indexed angle grid inevitably introduces approximation errors, causing the sampled contour points to deviate from the true boundary surface, which consequently affects the accuracy of vertebral shape reconstruction. Thus, to mitigate this deviation, a distance-based center should be marked by minimizing the average Euclidean distance between this center and the whole boundary surface. We formulate the distance-based center as spherical centroid $C_{s}$ based on Eq (6):
@@
where $\mathcal{F}(\cdot)$ represents the morphological dilation operation, $M$ refers to the segmentation mask of each vertebral volume, the boundary map $\mathcal{B}$ is the difference set between the dilated mask $\mathcal{F}(M)$ and the mask itself $M$. Besides, $d(\cdot)$ means the $L_{2}$ norm operation in Euclidean space. The first item in Eq (6) refers to the average distance from each boundary point $b$ to the center $c\in\mathcal{P}$. Concerning the non-convex shape prior of the spinous process, it will be more difficult to approximate finer contours of this region. Therefore, we attempt to push the spherical centroid closer to it. As illustrated by Fig.3, that will promote a denser sampling of contour points near the spinous process, then enhance the shape representations of the spinous process. That is why the second regularization term is incorporated, in which $\delta$ is the boundary point $j\in\mathcal{B}$ with the largest y-dimensional index $j_{y}$. $\lambda$ is a scaling parameter, empirically set as $0.005$.
@@ 3.3 Contour Generation with Explicit Contour Priors
Center decoder: Deep features from the encoder are transformed into regressed spherical centers via combinative operations of convolution, batch normalization, average pooling. The pooling process projects the feature dimension as 3, representing coordinates of the regressed center $\hat{C}=(\hat{c}_{x},\hat{c}_{y},\hat{c}_{z})$. The ground truth of center $C_{s}$ can be attained by Eq (6). And we employ the $L_{2}$ norm to regularize regressed centers $\hat{C}$, with the center loss $\mathcal{L}_{center}$ equal to $||\,C_{s}-\hat{C}\,||^{2}$. Concerning the limitation that some methods wrongly detect centers of specific pathological vertebrae, we aim to introduce Gaussian positional prompts as additional inputs for positional guidance. As shown in Fig.4(b), the positional prompt is defined as a Gaussian distribution $\mathcal{G}$ based on the following formula:
@@
Specifically, the mean value $\bm{\mu}=(\mu_{x},\mu_{y},\mu_{z})$ is set as the centroid based on the coarse prediction. For defining the variance $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$, we aim to cover the whole shape distribution using the given $\mathcal{G}$. Thus, as suggested by the two-sigma rule in mathematical statistics, the length of $4\bm{\sigma}$ should encompass the 3-dimensional length $\bm{m}=(m_{x},m_{y},m_{z})$ of each instance mask. Further, a threshold $\bm{\bar{m}}$ is introduced to ensure a wide enough region with attention weights decaying from the center $\bm{\mu}$ to the vertebral boundary. The center decoder will modify coarse centroids $\bm{\mu}$ as refined spherical centroids via the explicit constraint $\mathcal{L}_{center}$.
@@
$=\Large\tfrac{1}{|\mathcal{P}_{c}|}\displaystyle\sum_{p\in\mathcal{P}_{c}}\mathop{\min}\limits_{b\in\mathcal{B}}\sqrt{\|p-b\|^{2}}$ (10)
@@
Loss term: Guided by the recent work [38, 41], $\tfrac{1}{3}$ uniformly sampled boundary points can approximately restore the shape information of vertebrae. Thus, uniformly sampled boundary points from $\mathcal{B}$ is incorporated to calculate the contour loss. Besides, an additional binary mask supervision $\mathcal{L}_{mask}$ (cross-entropy + Dice loss) will exert implicit shape information. The binary mask decoder is configured with the structure of MedNeXt decoder [42], boosting the regression of vertebral contours. Thus, the total loss is defined as follows:
@@
3.4 Inference
In the inference phase, the patch to be refined in SLoRD is cropped along each vertebral center from coarse predictions by arbitrary segmentation networks as shown in Fig.4(a). The SLoRD framework will first refine three spherical centers corresponding to sequential vertebrae according to the prediction of the center decoder. Then the coefficient branch will output $3\times N$ (N refers to the dimension of $\bm{\rho}$) contour points surrounding three vertebral surfaces with different labels as revealed in Fig.5. After that, Marching Cube [46] is implemented to generate triangle meshes, which are further voxelized into a filled volume [47]. Lastly, we perform the binarized attention operation $\mathcal{O}_{B}$ on the set of coarse masks $\mathcal{S}_{C}$ and labeled masks $\mathcal{S}_{L}$ produced by SLoRD. The mathematical formulation is as follows:
$\mathcal{S}_{R}=\mathcal{O}_{B}(\mathcal{S}_{C},\mathcal{S}_{L})=\mathbb{1}(\mathcal{S}_{C})\cdot\mathcal{S}_{L}$ (12)
$\mathbb{1}(i)=\begin{cases}1,&\text{if }\;i>0\\ 0,&\text{otherwise }\end{cases}$ (13)
we can acquire the set of refined segmentation masks $\mathcal{S}_{R}$, which reveals both good label consistency and shape restoration via the auxiliary connected component analysis.